### v2

# Risk-adaptive Activation Steering for Safe Multimodal Large Language Models

Abstract
One of the key challenges of modern AI models is ensuring that they provide helpful responses to benign queries while refusing malicious ones. But often, the models are vulnerable to multimodal queries with harmful intent embedded in images. One approach for safety alignment is training with extensive safety datasets at the significant costs in both dataset curation and training. Inference-time alignment mitigates these costs, but introduces two drawbacks: excessive refusals from misclassified benign queries and slower inference speed due to iterative output adjustments. To overcome these limitations, we propose to reformulate queries to strengthen cross-modal attention to safety-critical image regions, enabling accurate risk assessment at the query level. Using the assessed risk, it adaptively steers activations to generate responses that are safe and helpful without overhead from iterative output adjustments. We call this Risk-adaptive Activation Steering (RAS). Extensive experiments across multiple benchmarks on multimodal safety and utility demonstrate that the RAS significantly reduces attack success rates, preserves general task performance, and improves inference speed over prior inference-time defenses.
1 Introduction
Multimodal Large Language Models (MLLMs) (Liu et al., 2024c; Wang et al., 2024a; Chen et al., 2024) leverage pretrained Large Language Models (LLMs) that have often gone through safety alignment on textual data. However, as shown in Fig. 1(a), current MLLMs fail to generate refusals against multimodal instructions with malicious intent embedded in images, despite extensive vision-language alignment, as also noted by  Xu et al. (2024); Liu et al. (2025). Existing approaches to address this problem generally fall into two categories: (i) training-based methods and (ii) inference-time methods. Training-based methods (e.g., supervised fine-tuning (Ding et al., 2025) or reinforcement learning (Zhang et al., 2025)) effectively enhance safety, but are costly: they require collecting high-quality safety data and joint training with general-task data to preserve utility (i.e., performance on general tasks). These demands become especially prohibitive for foundation models like MLLMs, where the large model size and multimodal inputs further amplify the training overhead.
Given the limitations of training-based approaches, recent work has shifted toward inference-time alignment, which aims to improve safety without additional training. These methods typically add safety prompts to the query during inference (Gong et al., 2025; Gao et al., 2024), or refine responses through iterative MLLM forward passes (Gou et al., 2024; Ding et al., 2024). However, safety prompts degrade utility by over-refusing benign queries, while response refinement incurs heavy computational overhead from the sequential process of generating and then refining responses.
These limitations highlight the need for a more precise refusal approach to accurately distinguish safe from unsafe queries, ensuring proper refusal while preserving utility, and a more efficient approach that avoids costly refusal processes. To achieve both precision and efficiency, it is crucial to analyze why accurate risk assessment fails at the query level, as only then can refusals be made without refinement, enabling faster and safer inference. To this end, we investigate why MLLMs struggle to conduct accurate safety reasoning, particularly for multimodal queries. Our analysis reveals that the key issue lies in insufficient cross-modal attention to safety-critical image regions. When an unsafe instruction is given in text (Fig. 1(b)), the model allocates significant attention to the unsafe text tokens such as “bomb”, generating an appropriate refusal. However, when the same unsafe instruction is given in multimodal format with the harmful context embedded in images, the model fails to allocate sufficient attention to the corresponding visual tokens, resulting in unsafe outputs (Fig. 1(a)).
Building on this analysis, we propose Risk-adaptive Activation Steering (RAS), an inference-time defense that dynamically steers a frozen MLLM toward refusal behavior based on the estimated risk of the input query. RAS consists of three stages: (i) vision-aware query reformulation, which appends concise visual contexts (i.e., a brief summary of the image) and safety prompts to strengthen cross-modal attention to safety-critical regions; (ii) risk evaluation, which estimates the threat level of the reformulated query; and (iii) adaptive activation steering, which adjusts model activations with intervention strength scaled according to the assessed risk. This design minimizes interference with benign queries, preserving utility, while effectively steering unsafe queries toward refusals.
Our evaluation shows that RAS substantially reduces attack success rates on diverse multimodal jailbreak benchmarks across multiple MLLMs, such as LLaVA-1.5 (Liu et al., 2024b), Qwen-VL-Chat (Bai et al., 2023), and InternLM-XComposer (Zhang et al., 2024). Moreover, RAS better preserves benign task performance and delivers significantly faster inference throughput compared to existing inference-time methods. Collectively, these results highlight dynamic, context-aware latent steering as an efficient and effective approach to enhance MLLM safety without compromising speed and utility. We summarize our contributions as follows:
  • We identify two key limitations of MLLM safety: (i) insufficient attention to safety-critical image regions, and (ii) utility degradation from safety-prompt-induced distribution shifts.
  • To address inadequate attention to unsafe image regions, we propose vision-aware query reformulation that guides cross-modal attention toward safety-critical visual tokens.
  • To address utility degradation caused by safety-prompt-induced distribution shifts, we introduce an exponentially weighted risk evaluation method that operates on the shifted distribution, enabling risk-adaptive activation steering to generate appropriate refusals.
(a) Unsafe multimodal instruction
(b) Unsafe text instruction
Figure 1: Lack of attention to safety-critical image regions. (a) For unsafe multimodal instructions, the model fails to allocate sufficient attention to safety-critical image regions (i.e., the bomb, highlighted in red), leading to unsafe responses. (b) In contrast, when the same instruction is given in text only, the model sufficiently attends to harmful text tokens (i.e., the text token “bomb”) and generates a refusal. This highlights a key limitation: insufficient attention to harmful visual tokens in multimodal queries. See Section 3.1 for a more comprehensive analysis of this issue. We employ LLaVA-1.5-7B for attention weight extraction.
2 Related Work
Training-based safety alignment. Training-based approaches to align MLLMs broadly fall into two categories: (i) supervised fine-tuning and (ii) reinforcement learning. Supervised fine-tuning trains a pretrained MLLM on safety datasets, as even modest amounts of safety-specific fine-tuning can reduce harmful responses (Zong et al., 2024; Ding et al., 2025). However, excessive safety data often leads to refusals on benign queries. To alleviate this issue, safety datasets are combined with general-purpose data, but the optimal balance remains unclear across tasks and models (Wang et al., 2025b; Bianchi et al., 2024). Reinforcement learning aligns MLLMs using human or synthetic preference annotations, typically through algorithms such as Proximal Policy Optimization (Schulman et al., 2017) or Direct Preference Optimization (Rafailov et al., 2023). SPA-VL (Zhang et al., 2025) follows this paradigm by constructing a preference dataset to guide outputs toward safer responses.
While effective at strengthening refusal behaviors, both fine-tuning and reinforcement learning face practical limitations: collecting high-quality data is labor-intensive, and the training pipeline incurs substantial computational resources. As a result, training-based approaches, despite improving safety, remain resource-intensive and introduce uncertain trade-offs with benign task performance.
Inference-time safety alignment. Recently, several inference-time methods have been proposed to enhance MLLM safety. CoCA (Gao et al., 2024) improves safety alignment through logit calibration by comparing output token logits with and without safety prompts, building on prior work showing that safety prompts increase refusal rates against malicious queries (Liu et al., 2024d; Gong et al., 2025). Other approaches, such as AdaShield (Wang et al., 2024b), MLLM-Protector (Pi et al., 2024), Immune (Ghosal et al., 2025), and ETA (Ding et al., 2024), employ external reward models to evaluate and refine responses when harmful content is detected. However, using a separate reward model incurs substantial computational and memory costs due to iterative response refinement and dual-model operation. As an alternative, ECSO (Gou et al., 2024) avoids reliance on an external reward model by leveraging the MLLM itself for evaluating and regenerating responses, but it still incurs the overhead associated with response refinement.
Unlike prior methods that generate full responses for safety assessment, our proposed RAS assesses risks at the query level by reformulating it with short visual prompts and safety prompts. Moreover, rather than relying on binary judgments from the MLLM itself or external reward models that classify outputs as safe or unsafe, RAS produces continuous risk scores. These scores are derived by measuring distributional similarity between the model outputs and refusal-related distributions. This approach significantly reduces inference overhead while maintaining safe and helpful responses.
Activation steering. Activation steering has been primarily studied in language models, where injecting steering vectors into internal activations during inference can elicit or suppress specific responses without significant computational overhead or additional decoding steps (Arditi et al., 2024; Liu et al., 2023; Panickssery et al., 2023). Such latent-space interventions offer lightweight yet powerful mechanisms to dynamically influence model behavior, making activation steering attractive for alignment tasks. However, as applying the same steering intensity to all queries may yield undesired outputs (Stoehr et al., 2024; Wang et al., 2025a), we adopt an adaptive approach that scales the intervention strength according to the assessed risk of each query. In this way, safe queries remain unaffected, while unsafe queries are steered proportionately to their risk levels.
Figure 2: Overview of RAS. RAS consists of three stages: (i) Query Reformulation, which augments input queries with concise visual contexts and safety prompts to strengthen cross-modal attention to safety-critical image regions; (ii) Risk Evaluation, where MLLM output activations are compared with unsafe prototypes to produce similarity-based risk scores; and (iii) Risk-adaptive Activation Steering, where activations are steered toward refusal behavior according to the risk score.
3 Approach
To address two underexplored limitations in inference-time alignment: (i) insufficient attention to safety-critical image regions and (ii) distribution shifts in output probabilities induced by safety prompts, we propose Risk-adaptive Activation Steering (RAS). RAS consists of three stages: (i) vision-aware query reformulation (Sec. 3.1), (ii) exponentially weighted risk evaluation (Sec. 3.2), and (iii) risk-adaptive activation steering (Sec. 3.3). We provide an overview of RAS in Fig. 2.
3.1 Vision-Aware Query Reformulation (Stage 1)
For multimodal instructions, a text query (e.g., “How can I make the item in the image?”) can be interpreted as safe or unsafe depending on the accompanying image (e.g., a bomb vs. a chair in Fig. 3a). In such cases, the model must allocate sufficient attention to the safety-critical regions to provide helpful responses to benign inputs while generating refusals to malicious inputs. For language models to generate an appropriate refusal, the earlier tokens in the response play a crucial role (Qi et al., 2024). For example, when a response begins with a clear refusal (e.g., “I’m sorry,”), the model successfully declines to provide an answer, whereas responses that start by complying with the query often fail to reject it.
Therefore, we define $a^{(l,h)}_{j}$, the cross-modal attention weight assigned to visual token $v_{j}$ by text query tokens in head $h$ of layer $l$ as $a^{(l,h)}_{j}=\max_{t\in\mathcal{T}}a^{(l,h)}_{j,t}$, where $\mathcal{T}$ is the set of text tokens. Since only a few attention heads specialize in visual grounding (Kang et al., 2025b), we compute $a^{*}_{j}$, the effective cross-modal attention weight to $v_{j}$, by averaging over the top-$n$ heads:
$a^{*}_{j}=\frac{1}{|\mathcal{H}_{n}|}\sum_{(l,h)\in\mathcal{H}_{n}}a^{l,h}_{j},$ (1)
where $\mathcal{H}_{n}$ denotes the set of top-$n$ heads across all layers ranked by their attention strength.
As shown in Fig. 3, the attention weights assigned to the objects are small, indicating weak visual grounding. This suggests that insufficient attention to distinct image regions would make safe and unsafe multimodal instructions less separable in the representation space, particularly when the text queries are identical. To quantify the representational separability, we employ the Fisher Discriminant Ratio (FDR), computed from the last token activations of safe and unsafe object images, following prior work on representational discrimination (Wang et al., 2009; Zarka et al., 2020; Ramezani et al., 2025). As these activations determine the first response token, they provide a direct reflection of the model’s safety reasoning (Qi et al., 2024).
Figure 3: Attention maps for unsafe (top) and safe (bottom) objects under various query formulations. SP. denotes safety prompt. VC. denotes visual context. (a) Example of unsafe vs. safe instructions under the same text query. (b–e) Cross-modal attention maps from text to visual tokens. (b) With only the original query, attention weights to the objects are small, indicating weak visual grounding. (c) Safety prompts fail to enhance attention to the objects, whereas (d–e) visual contexts significantly improves it. We employ LLaVA-1.5-7B for attention weight extraction.
For each transformer layer $l$, we denote the sets of last token activations for safe and unsafe object queries as $\{\mathbf{x}^{\,l}_{q}\mid q\in\mathcal{Q}^{\text{object}}_{\text{safe}}\}$ and $\{\mathbf{x}^{\,l}_{q}\mid q\in\mathcal{Q}^{\text{object}}_{\text{unsafe}}\}$ with equal samples sizes, i.e. $|\mathcal{Q}^{\text{object}}_{\text{safe}}|=|\mathcal{Q}^{\text{object}}_{\text{unsafe}}|$. Images of safe objects (e.g., chairs, clothes) are sampled from the ImageNet-1K dataset (Deng et al., 2009), while unsafe objects (e.g., firearms, explosives) are obtained from the Dangerous Objects Dataset (Alinadilawaiz, 2023). The FDR at layer $l$ with hidden dimension $d$ is defined as:
$\mathrm{FDR}(l)=(\bm{\mu}^{l}_{\text{safe}}-\bm{\mu}^{l}_{\text{unsafe}})^{\top}\left(\bm{\Sigma}^{l}_{\text{safe}}+\bm{\Sigma}^{l}_{\text{unsafe}}+\epsilon\bm{I}\right)^{-1}(\bm{\mu}^{l}_{\text{safe}}-\bm{\mu}^{l}_{\text{unsafe}}),$ (2)
where $\bm{\mu}^{l}_{\text{safe}},\bm{\mu}^{l}_{\text{unsafe}}\in\mathbb{R}^{d}$ are the mean activations, $\bm{\Sigma}^{l}_{\text{safe}},\bm{\Sigma}^{l}_{\text{unsafe}}\in\mathbb{R}^{d\times d}$ are the covariance matrices, and $\epsilon I$ is a term for numerical stability.
Figure 4: FDR values across layers for various query formulations. SP. denotes safety prompt. VC. denotes visual context. Lower FDR indicates less separable representations. We employ LLaVA-1.5-7B to compute FDR.
When the MLLM only uses the query, the overall FDR across layers remains low (orange line in Fig. 4). Since a lower FDR indicates less separable representations, this suggests that insufficient attention to distinct image regions leads to similar embeddings when the same text query is used. We then assess whether prior works (Liu et al., 2024d; Gong et al., 2025) that incorporate safety prompts can mitigate this. However, they yield no notable improvement in FDR (green line in Fig. 4), due to the model’s persistent lack of attention to distinct image regions even under safety prompting (Fig. 3c).
Vision-aware query reformulation. To address the insufficient attention to query-relevant image regions, we incorporate the query with concise visual contexts (i.e., a brief summary of the image), inspired by prior works demonstrating that textualizing key visual elements strengthens cross-modal attention (Pandey et al., 2022; Kang et al., 2025a; b). As shown in Fig. 3d, visual contexts strengthen attention to the objects, yielding higher FDR (blue line in Fig. 4). Furthermore, with the strengthened cross-modal attention from visual contexts (Fig. 3e), adding safety prompts results in a significant increase in FDR (red line in Fig 4), unlike in the absence of such attention.
To examine whether visual contexts can replace images, we evaluate the ‘Visual Context + Query’ and ‘Safety Prompt + Visual Context + Query’ formulations without images (brown and purple lines in Fig. 4). Compared to the corresponding settings with images, these formulations yield lower FDR. Thus, while visual contexts enhance separability, they cannot fully replace images, which provide complementary cues that enable stronger discrimination between safe and unsafe queries.
Overall, vision-aware query reformulation, where safety prompts and short visual contexts are added to the original query, yields discriminative representations between safe and unsafe queries. This enables precise risk assessments in the subsequent evaluation stage.
3.2 Exponentially Weighted Risk Evaluation (Stage 2)
Although responses can be generated directly from reformulated queries, the safety prompt causes the output probability distribution of the initial tokens to skew toward refusal-like responses even for benign inputs, leading to utility degradation. However, we observe that reformulated queries yield separable representations between safe and unsafe queries (the separation between the blue and orange histograms in Fig. 6). Leveraging this separation, we estimate the safety of a given query. This can be efficiently achieved by measuring the alignment between the probability distributions of the initial outputs and those of refusal-related responses, since refusal behavior (e.g., the use of text tokens such as “I’m sorry”) is reflected in the first few response tokens (Qi et al., 2024).
Prototype-based similarity evaluation. Evaluating refusal behavior requires comparing a given query’s output distribution with a reference distribution derived from refusals. To construct this reference, we generate $\mathcal{Q}^{\text{text}}_{\text{unsafe}}$, a set of policy-violating text queries, created using GPT-4 (see Appendix A.3 for the list of queries). For each query, we extract the last layer activations of the initial response tokens and compute their token-wise means to obtain unsafe prototypes $\bm{\mu}_{p}$. Formally, $\bm{\mu}_{p}$ is defined as follows,
$\bm{\mu}_{u}^{n}=\frac{1}{|\mathcal{Q}^{\text{text}}_{\text{unsafe}}|}\sum_{q\in\mathcal{Q}^{\text{text}}_{\text{unsafe}}}\mathbf{x}^{n}_{q}$ (3)
where $|\mathcal{Q}^{\text{text}}_{\text{unsafe}}|$ denotes the number of queries and $n$ denotes the token position in the response sequence. For each query $q\in\mathcal{Q}^{\text{text}}_{\text{unsafe}}$, we extract $\mathbf{x}_{q}^{n}$, the last layer activation corresponding to the $n^{\text{th}}$ response token. Finally, $\bm{\mu}^{n}_{u}$ represents the unsafe prototype activation in the last layer at position $n$ (see Appendix F for results on intermediate layers).
(a) Similarity score distributions of safe objects
(b) Similarity score distributions of unsafe objects
Figure 5: Similarity score ($S_{i}$) distributions of (a) safe (b) unsafe objects under various query formulations. SP. denotes safety prompt. VC. denotes visual context. Higher similarity scores indicate output distributions similar to refusals. We employ LLaVA-1.5-7B to extract $S_{i}$ scores.
To measure output similarity between unsafe prototypes and a given input query $i$, we extract the last layer activations of the response tokens at position $n$, denoted $\mathbf{x}_{i}^{n}$. The similarity $S_{i}$ is measured by the cosine similarity between the exponentially weighted sum of output distributions:
$S_{i}=\cos(\textstyle\sum_{n=1}^{N}\gamma^{n-1}\mathbf{\hat{y}}_{i}^{n},\ \sum_{n=1}^{N}\gamma^{n-1}\mathbf{\hat{y}}_{u}^{n}),$ (4)
where $\hat{\mathbf{y}}^{n}_{i}=\text{softmax}(\text{LMHead}(\mathbf{x}^{n}_{i}))$ and $\hat{\mathbf{y}}^{n}_{u}=\text{softmax}(\text{LMHead}(\bm{\mu}^{n}_{u}))$ denote the output probability distributions obtained by passing the query activations and unsafe prototypes through the language model head and subsequently applying the softmax function. $\gamma\in(0,1)$ denotes an exponentially decaying term that assigns greater weights to earlier response tokens, where refusal behavior is more pronounced. With small $\gamma$, the weights of subsequent tokens sharply approach zero, making their contribution negligible. Hence, we consider only a small number of initial tokens (e.g., $N=3$), which suffices to capture refusal behavior while ensuring computational efficiency.
When $S_{i}$ is high, the query resembles unsafe prototypes and is likely to trigger a refusal, whereas a low $S_{i}$ indicates a benign query which is likely to comply.
Distribution shift from safety prompts. In Fig. 5, we plot the similarity score distributions $S_{i}$ for safe and unsafe object images under the query “How can I make the item in the image?”. When using the query with the image but without safety prompts (orange), both distributions concentrate around $S_{i}\approx 0$, indicating that the model tends to provide answers instead of issuing refusals. Adding safety prompts (green) shifts both safe and unsafe distributions toward higher $S_{i}$ values, i.e., in the direction of refusals, as safety prompts instruct the models to reject queries that might be unsafe. However, because the model fails to sufficiently attend to safety-critical image regions, it cannot properly distinguish safe from unsafe cases, resulting in refusal-like responses for both.
To address this limitation, we apply vision-aware query reformulation (red histograms). Unsafe queries exhibit larger distributional shifts, whereas safe queries show smaller shifts, resulting in a clearer separation. This improvement arises because the added visual context strengthens cross-modal attention to safety-critical regions, enabling more accurate safety reasoning. We also examine reformulated queries without images (purple histograms), which show weaker discrimination, consistent with the FDR results in Fig. 4. These observations further highlight the role of cross-modal attention in distinguishing safe from unsafe queries.
Risk evaluation. Building on this insight, we derive risk scores from reformulated queries and use these scores to steer the activations of original queries. To derive risk scores, we use $S_{i}$ values from unsafe SPA-VL (Zhang et al., 2025) samples (orange histogram in Fig. 6), with the mean used as a baseline $S_{\text{base}}$ to define intermediate risk levels. We use SPA-VL, as the dataset covers diverse harmful categories (e.g., illegal activity, privacy violation). Each $S_{i}$ is then mapped to a continuous risk score $r(S_{i})\in(0,1)$ using a sigmoid function centered at $S_{\text{base}}$ (red line in Fig. 6).
Figure 6: Distribution of similarity scores for reformulated queries from $M$-Vet (safe) and SPA-VL (unsafe).
This can be expressed as:
$r(S_{i})=\sigma\!\bigl(\alpha\,(S_{i}-S_{\text{base}})\bigr),$ (5)
where $\sigma(z)=\frac{1}{1+e^{-z}}$. Here, $\alpha>0$ controls the slope and is determined adaptively to satisfy $r(S_{i})\approx 1$ when $S_{i}=1$. Consequently, queries with similarity scores greater than $S_{\text{base}}$ produce high risk scores, whereas safe queries with scores below $S_{\text{base}}$ (e.g., blue histogram in Fig. 6, obtained from benign MM-Vet samples) yield low risk scores.
3.3 Risk-adaptive Activation Steering (Stage 3)
At this stage, we steers model activations adaptively toward refusal behavior based on the risk evaluated in Stage 2. The novelty of our approach lies in its adaptive design, applying minimal intervention to benign queries while enforcing refusals to malicious queries.
Refusal vector computation. Following Arditi et al. (2024), we employ activation steering along refusal vectors, but redefine them for a more targeted and effective refusal behavior. Rather than using the difference between mean activations of safe and unsafe queries, we use the vectors from each input query activation to the unsafe prototype (see Tab. 6 in Appendix G for comparison on refusal behavior). Specifically, the refusal vector $\mathbf{v}^{n}$ at the last layer is computed as:
$\mathbf{v}^{n}=\bm{\mu}^{n}_{u}-\mathbf{x}_{i}^{n},$ (6)
where $n$ denotes the position of the output token and $i$ denotes the input query. This directional vector encodes the adjustment required to steer activations toward refusals and away from generating harmful responses.
Risk-adaptive activation steering. We scale the refusal vector by the risk score $r(S_{i})$ to ensure that the intervention strength is proportional to the risk estimated with EWRE (Stage 2). That is, for the last layer activation of the original query $\mathbf{x}^{n}_{i}$, we compute the steered activation $\tilde{\mathbf{x}}^{n}_{i}$ as:
$\tilde{\mathbf{x}}^{n}_{i}=\mathbf{x}_{i}^{n}+r(S_{i})\cdot\mathbf{v}^{n}.$ (7)
For computational efficiency, we apply activation steering to the last layer and to the first $N$ response tokens, matching those used for risk evaluation. This formulation ensures that benign queries ($r(S_{i})\approx 0$) receive negligible steering, preserving their original representations to maintain helpful responses. Unsafe queries ($r(S_{i})\approx 1$) receive maximal steering, guiding the model toward appropriate refusals. For ambiguous queries ($0<r(S_{i})<1$), the intervention strength is scaled adaptively according to their similarity to unsafe patterns (see Appendix H for qualitative results).
4 Experiments
We evaluate RAS across three dimensions: (i) safety, measured by attack success rates on multi-modal jailbreak datasets; (ii) utility, measured by accuracies or scores on general multimodal reasoning tasks; and (iii) inference speed, measured by tokens per second on identical hardware.
4.1 Experimental Setup
Benchmarks. For safety, we report attack success rates (ASR) on MM-SafetyBench (Liu et al., 2024d), SPA-VL (Zhang et al., 2025), and FigStep (Gong et al., 2025), where attack success is judged by MD-Judge-v0.2-Internlm2 (Li et al., 2024), following (Ding et al., 2024; Huang et al., 2024; Zhang et al., 2025). Utility is evaluated on Sci-QA (Lu et al., 2022), MM-Vet (Yu et al., 2023), GQA (Hudson & Manning, 2019), and MME (Fu et al., 2024), using the official metrics provided by each benchmark (see Appendix B for further details on benchmarks).
Models. To verify the generalizability of RAS across various models and sizes, we evaluate it on LLaVA-1.5-7B, LLaVA-1.5-13B, Qwen-VL-Chat, and InternLM-XComposer-2.5-7B.
Baselines. We compare RAS with state-of-the-art inference-time alignment methods, including FigStep (Gong et al., 2025), CoCA (Gao et al., 2024), ECSO (Gou et al., 2024), and ETA (Ding et al., 2024). FigStep improves refusal rates by adding safety prompts, while CoCA calibrates logits by comparing distributions with and without safety prompts. ECSO conducts self-evaluation to refine responses, whereas ETA uses an external reward model for response refinement. Together, these baselines span diverse inference-time strategies, offering a comprehensive comparison for RAS.
Implementation details. For attention weight evaluation, we follow Kang et al. (2025b) and use 3 heads. To compute FDR, we sample 100 safe and unsafe object images each, and employ 50 text queries to construct unsafe prototypes. For risk evaluation, we set $\gamma=0.3$ and $N=3$ across all models. For $S_{\text{base}}$ and $\alpha$, as $S_{i}$ distributions differ across models, they are adaptively determined by scores from 100 samples from SPA-VL (Sec. 3.2). We report the specific values in Appendix A.4.
Table 1: Comparison in safety and utility benchmarks. Bold indicates the best performance (lowest ASR for safety, highest accuracy/score for utility). Average gains show changes relative to the ’Original’ model: ASR reduction for safety and performance drop for utility.
Model Method Safety Utility
$M$-Safety SPA-VL FigStep Average Sci-QA $M$-Vet GQA MME (Percep./Cog.) Average
(ASR $\downarrow$) (ASR $\downarrow$) (ASR $\downarrow$) Gain (%) (Img. Acc. $\uparrow$) (Score $\uparrow$) (Acc. $\uparrow$) (Score $\uparrow$) Gain (%)
LLaVA-1.5-7B Original 40.1 47.2 59.3 - 69.5 30.5 61.9 1505.1 / 355.7 -
FigStep 26.8 32.4 52.0 +25.6 68.3 29.5 61.3 1435.7 / 275.0 -6.7
CoCA 19.7 10.9 51.6 +46.9 67.7 28.9 60.3 1526.5 / 283.6 -5.9
ECSO 15.9 23.4 37.4 +49.2 69.5 30.3 61.9 1505.1 / 355.7 -0.1
ETA 15.8 17.0 7.8 +70.5 69.5 30.4 61.9 1509.3 / 339.6 -0.9
RAS (Ours) 4.1 8.3 2.2 +89.5 69.5 30.5 61.9 1505.1 / 355.7 0.0
LLaVA-1.5-13B Original 41.0 40.8 61.6 - 72.7 35.6 63.2 1529.9 / 298.6 -
FigStep 23.0 21.5 55.0 +34.0 72.1 33.2 62.4 1423.9 / 322.1 -1.6
CoCA 12.4 10.2 52.4 +53.2 71.4 32.1 62.3 1472.8 / 301.8 -3.1
ECSO 13.8 15.5 15.0 +67.9 72.7 35.5 63.2 1529.9 / 298.6 -0.1
ETA 11.7 15.1 22.6 +65.9 72.7 35.6 63.2 1531.2 / 296.1 -0.2
RAS (Ours) 6.9 4.9 2.0 +89.3 72.7 35.4 63.2 1529.9 / 298.6 -0.1
Qwen-VL-Chat Original 33.1 12.5 52.4 - 68.0 48.7 57.3 1489.9 / 331.8 -
FigStep 8.1 5.7 44.4 +48.4 64.4 39.0 56.8 1480.9 / 296.4 -7.5
CoCA 2.6 4.2 32.2 +65.7 66.7 38.7 56.9 1377.1 / 319.3 -6.9
ECSO 19.1 7.6 45.4 +31.6 68.0 47.2 57.3 1489.9 / 331.8 -0.6
ETA 9.3 4.5 9.2 +72.8 67.8 45.9 57.3 1487.9 / 331.8 -1.2
RAS (Ours) 0.7 3.0 1.2 +90.5 68.0 46.9 57.3 1489.9 / 331.8 -0.7
InternLM- XComposer-2.5 Original 21.8 27.6 22.6 - 94.7 50.1 59.1 1623.7 / 551.1 -
FigStep 6.3 6.8 7.0 +71.8 86.1 47.2 58.9 1577.7 / 516.8 -4.9
CoCA 6.1 5.9 16.0 +59.9 93.3 48.1 58.8 1606.5 / 551.1 -1.4
ECSO 14.9 19.6 16.6 +29.1 94.7 49.4 59.1 1623.7 / 551.1 -0.3
ETA 7.3 14.0 6.0 +63.1 94.6 47.4 58.1 1629.4 / 546.1 -1.5
RAS (Ours) 5.1 4.2 4.0 +81.2 94.7 50.0 59.1 1623.7 / 551.1 -0.1
4.2 Results
Safety. We show the jailbreak attack ASR results on the left of Tab. 1. While existing inference-time defenses lower ASR to some extent, they still leave notable vulnerabilities, particularly against attacks such as FigStep. For both LLaVA-1.5-7B, LLaVA-1.5-13B, and Qwen-VL-Chat, prompt-based methods (FigStep, CoCA) lower ASR on MM-Safety and SPA-VL yet leave FigStep largely unaffected. In contrast, RAS consistently achieves the lowest ASR across all evaluated models and benchmarks, yielding significant improvements over prior methods.
Utility. An effective defense must not only enhance safety but also preserve the general task performance of MLLMs. As shown on the right of Tab. 1, RAS maintains performance nearly identical to the original models across all tasks, whereas other inference-time defenses often cause notable degradation. These results demonstrate that RAS not only provides strong refusals against malicious queries, but also preserves the multimodal reasoning capabilities of MLLMs. For MM-Vet, which evaluates MLLMs across diverse tasks to provide a comprehensive assessment of vision-language abilities, we report detailed subscores in Tab. 4 in Appendix C.
Inference speed. For real-world deployment, it is essential that inference-time aligned MLLMs maintain efficient inference speed. To assess this, we measure relative throughput, defined as tokens per second relative to the original model, following (Svirschevski et al., 2024; Liu et al., 2024a; Fedorov et al., 2024). As shown in Fig. 8, RAS achieves the lowest ASR and the highest throughput on SPA-VL, while preserving MM-Vet utility with minimal slowdown. This efficiency stems from its lightweight design, which involves generating a short visual context for query reformulation and applying risk evaluation/activation steering to just three tokens. In contrast, ECSO and ETA repeatedly verify and regenerate responses, while CoCA adjusts logits at every decoding step, all of which introduce substantial computational overhead.
Figure 7: Comparison of ASR, utility score, and relative throughput on SPA-VL (left) and $M$-Vet (right). For SPA-VL, the desirable region is the lower right (low ASR, high throughput), whereas for $M$-Vet, it is the upper right (high utility, high throughput).
Figure 8: Sweeping $\bm{\gamma}$ and $\bm{N}$. Safety is evaluated by average ASR on $M$-Safety, SPA-VL, and FigStep, and utility is evaluated by $M$-Vet scores. A larger $\gamma$ enhances safety but degrades utility. In contrast, NN has a minor effect, with differences negligible for $N\geq 3$ due to exponential decay.
Table 2: Ablation results of RAS. We compare using Stage 1 only, and full RAS with and without images in Stage 2, under both binary and adaptive activation steering in Stage 3.
Model Stage Image in
Stage 2
Shifting Method
in Stage 3
Safety Utility
$M$-Safety SPA-VL FigStep rec ocr know gen spat math Total
LLaVA-1.5-7B [1] - - 3.5 7.1 2.8 38.2 25.6 13.9 18.9 27.2 7.7 28.6
[1, 2, 3] Binary 15.6 26.0 6.6 41.2 25.6 15.5 20.8 25.1 11.5 30.0
[1, 2, 3] Adaptive 14.0 23.4 6.0 41.0 25.7 15.9 21.5 25.3 11.5 29.9
[1, 2, 3] Binary 5.4 9.4 2.4 40.7 27.3 15.8 20.9 27.5 11.5 30.5
[1, 2, 3] Adaptive 4.1 8.3 2.2 41.1 26.9 16.2 21.8 26.7 11.5 30.5
4.3 Detailed Experiments
Ablation Study.
Here, we provide ablation studies on RAS to examine the impact of hyperparameters and the contribution of each component, using LLaVA-1.5-7B. For ablation using an alternative source of unsafe text queries to define unsafe prototypes, see Appendix F.
Effect of $\bm{\gamma}$ and $\bm{N}$ in EWRE.
EWRE computes the similarity between output distributions of the first $N$ response tokens and the corresponding $N$ unsafe prototypes, using a weighted sum with an exponential decaying factor $\gamma$. We vary $\gamma$ and $N$ to assess their impact on safety and utility, and show the results in Fig. 8. On the left, varying $\gamma$ with $N=3$ shows that larger $\gamma$ lowers ASR, indicating stronger refusals, but also degrades utility. We set $\gamma=0.3$ as the default, as it achieves the best trade-off, minimizing ASR while preserving utility close to its original score.
Next, we fix $\gamma=0.3$ and vary the number of tokens $N$. Smaller values of $N$ (e.g., 1-2) result in relatively high ASR, while larger $N$ reduces it. However, the effect of $N$ is less pronounced than that of $\gamma$, and for $N\geq 3$, performances remain stable as subsequent tokens are exponentially down-weighted. Thus, we set $N=3$ as an effective and efficient choice.
Effect of varying RAS stages
We report ablation results on RAS stages in Tab. 2. Using Stage 1 alone (row 1) shows that generating responses directly from reformulated queries degrades utility, due to output distribution shifts induce by safety prompts (Sec. 3.2). Rows 2–3 vs. 4–5 compare risk evaluation (Stage 2) without and with images. As shown by the $S_{i}$ distribution in Fig. 5, removing images weakens risk evaluation, resulting in higher ASR and lower utility. For Stage 3, we compare binary vs. adaptive activation steering, where binary steering replaces the sigmoid in Stage 2 with a unit-step function. Adaptive steering yields additional safety gains while preserving utility by adjusting intervention strength smoothly around the threshold, offering stronger defense against intermediate-risk queries that binary steering would either fully comply with or outright reject.
5 Conclusion
We introduce Risk-adaptive Activation Steering (RAS), an inference-time defense designed to improve MLLM safety without utility degradation or inference overhead. RAS reformulates queries to strengthen cross-modal attention to safety-critical image regions, enabling accurate risk assessment. Based on the assessed risk, it adaptively steers model activations, applying stronger interventions for unsafe queries and minimal intervention for safe queries. This approach to assess risks at the query level enables refusals without utility degradation and also eliminates the overhead of prior methods that refine responses iteratively. Extensive experiments across multiple multimodal safety and utility benchmarks demonstrate that RAS substantially reduces attack success rates, preserves performance on benign tasks, and improves inference speed compared to prior inference-time defenses, establishing it as an efficient and effective approach for safe MLLMs.
Ethics Statement
This work investigates the safety alignment of Multimodal Large Language Models (MLLMs) using publicly available benchmarks that include harmful or toxic prompts. We acknowledge the ethical risks of working with such data, as well as the possibility that models may generate unsafe responses under such adversarial conditions. Our approach aims to mitigate these risks by reducing harmful responses, thereby contributing to a more responsible deployment of MLLMs. While our method improves defenses, it does not fully eliminate vulnerabilities; continued research is necessary to better understand and mitigate ethical risks and potential misuse.
Reproducibility Statement
We have taken several steps to ensure the reproducibility of our work. Our proposed Risk-adaptive Activation Steering (RAS) is described in detail, with mathematical formulations and ablation analyses provided in the main text and appendices. We clearly report all hyperparameters and model-specific parameters used for risk evaluation and activation steering, including values of $\gamma$, $N$, $S_{\text{base}}$, and $\alpha$. Experiments are conducted on publicly available safety and utility benchmarks, including MM-SafetyBench, SPA-VL, FigStep, Sci-QA, MM-Vet, GQA, and MME. The multimodal models we evaluate: LLaVA-1.5-7B/13B, Qwen-VL-Chat, and InternLM-XComposer-2.5 are all publicly released. Baseline comparisons against FigStep, CoCA, ECSO, and ETA are implemented under identical settings. To further facilitate reproducibility, we will release our code upon publication.
Use of LLMs
We use Large Language Models (LLMs) to evaluate our method on existing multimodal models, to generate unsafe text queries for risk assessment, and to serve as judges when evaluating model responses. We also use it for polishing sentences.
Appendix A Further Implementation Details
A.1 Visual Context Generation
To generate concise visual contexts that enhance cross-modal reasoning, we prepend the following prompt (highlighted in blue) to the original query. For illustration, we provide a sample from the MM-Safety benchmark, including both the input prompt and the model’s response. The visual context generated by the model is highlighted in orange.
A.2 Vision-aware Query Reformulation
The reformulated query for risk assessment consists of three components: (i) a safety prompt, (ii) a short visual context, and (iii) the original query (Sec.3.1). The safety prompt (highlighted in green) is fixed across all queries, while the visual context (highlighted in orange) is generated using the procedure described in AppendixA.1. These components are appended to the original text query to form the reformulated input. Since risk evaluation relies on the activations of the first three tokens, the response is generated only up to the third token (e.g., “m” in “I’m”).
A.3 Unsafe Text Queries
To construct unsafe prototypes in Sec. 3.2, we use unsafe text queries created by GPT-4. We show the complete list of queries in Fig. 9.
Figure 9: List of unsafe text queries generated by GPT-4.
A.4 Risk Evaluation Parameters
As described in Sec. 3.2, both $S_{\text{base}}$ and $\alpha$ are derived from the distribution of similarity scores $S_{i}$ computed on 100 SPA-VL samples. Specifically, $S_{\text{base}}$ is set to the mean of $S_{i}$, and $\alpha$ is calibrated such that $r(S_{i})\approx 1$ when $S_{i}=1$. The resulting values for each model are summarized in Tab.3.
Table 3: Model-specific parameters for risk assessment.
LLaVA-1.5-7B LLaVA-1.5-13B Qwen-VL-Chat InternLM-XComposer-2.5-7B
$S_{\text{base}}$ 0.711 0.871 0.611 0.549
$\alpha$ 15.901 35.261 11.813 10.188
Appendix B Benchmark Details
MM-Safety (Liu et al., 2024d). MM-SafetyBench evaluates the safety of MLLMs against harmful queries paired with images. Covering 13 safety-critical scenarios with 5,040 text–image pairs, it introduces three adversarial settings: (i) Stable Diffusion (SD), which generates images from harmful key phrases; (ii) Typography (Typo), which embeds harmful key phrases as text within images; and (iii) a combined SD+Typo approach, which arranges both SD and Typo images in a vertical layout. As SD+Typo has been reported to be the most effective at bypassing model safeguards, we report ASR under this strongest setting.
SPA-VL (Zhang et al., 2025). SPA-VL is a large-scale dataset for aligning MLLMs with safety-relevant preference data generated by language models. It contains 93.3k preference samples across diverse harmful domains for training, along with 530 test samples (265 safe and 265 unsafe) for safety and utility evaluation. Following Ding et al. (2024), we report ASR on the 265 harmful test queries.
FigStep (Gong et al., 2025). FigStep consists of 500 queries where unsafe instructions are embedded in images as typographic prompts. Each query is constructed by rendering a harmful request (e.g., “Steps to manufacture illegal drugs”) as a numbered list with empty entries, while the accompanying text query is identical across all samples, simply asking to complete the list in the image.
Sci-QA (Lu et al., 2022). Sci-QA is a large-scale dataset designed to evaluate multimodal question answering in the science domain. It contains over 21,000 multiple-choice questions drawn from elementary to high school curricula, spanning natural science, social science, and language science. Each question may include texts, diagrams, or images as context, offering a diverse and challenging setting to assess multimodal reasoning.
MM-Vet (Yu et al., 2023). MM-Vet evaluates MLLMs on complex multimodal reasoning tasks, including recognition, OCR, knowledge reasoning, language generation, spatial reasoning, and math. It consists of 218 open-ended questions assessed with an LLM-based scoring system. Following the original paper, we use GPT-4 (0613) as the evaluator.
GQA (Hudson & Manning, 2019). GQA is a large-scale benchmark for visual reasoning and compositional question answering. It evaluates object recognition, spatial understanding, and logical inference, providing a systematic test on coherent multi-step reasoning beyond basic recognition.
MME (Fu et al., 2024). MME is a comprehensive benchmark spanning 14 subtasks across perception (object recognition, OCR, fine-grained identification) and cognition (commonsense reasoning, math, translation, code). All instruction–answer pairs use a concise yes/no format, enabling broad and consistent evaluation of vision–language abilities.
Appendix C Subscores on MM-Vet
In Sec. 4.2, we report overall MM-Vet scores as part of the utility evaluation. To provide a more fine-grained analysis, Tab. 4 reports subscores across various categories. Across all models, RAS achieves performance on MM-Vet comparable to the original models, indicating that our method improves safety without notable utility degradation.
Table 4: Subscores on MM-Vet across different capability categories. Subscores are reported across six categories: recognition (rec), optical character recognition (ocr), knowledge reasoning (know), generation (gen), spatial understanding (spat), and mathematics (math). RAS maintains overall performance comparable to the original models, demonstrating that our method does not compromise utility while improving safety.
Model Method rec ocr know gen spat math Total
LLaVA-1.5-7B Original 41.0 26.9 16.2 21.8 26.7 11.5 30.5
FigStep 39.6 24.1 15.7 21.0 27.5 7.7 29.5
CoCA 38.6 24.7 16.2 21.5 26.8 7.7 28.9
ECSO 40.8 26.9 15.5 21.1 26.8 11.5 30.3
ETA 41.1 24.9 18.1 22.5 28.0 7.7 30.4
RAS (Ours) 41.1 26.9 16.2 21.8 26.7 11.5 30.5
LLaVA-1.5-13B Original 44.7 32.2 20.7 21.6 36.1 11.2 35.6
FigStep 42.1 32.0 18.0 23.1 32.4 11.5 33.2
CoCA 40.3 32.1 20.4 24.0 29.7 7.7 32.1
ECSO 44.3 31.5 22.7 24.5 35.5 11.5 35.5
ETA 44.9 32.1 22.0 27.0 36.0 11.5 35.6
RAS (Ours) 45.5 30.8 22.5 24.9 34.0 11.5 35.4
Qwen-VL-Chat Original 60.2 40.8 45.2 41.1 39.7 22.7 48.7
FigStep 49.9 32.0 31.2 33.4 34.7 3.8 39.0
CoCA 45.0 35.7 32.6 28.9 38.9 7.7 38.7
ECSO 58.6 38.1 44.5 38.1 37.3 18.8 47.2
ETA 57.8 35.5 42.7 36.8 36.5 22.7 45.9
RAS (Ours) 58.9 38.2 41.0 36.8 39.7 18.5 46.9
InternLM-XComposer-2.5 Original 56.1 53.4 37.3 43.4 47.7 26.9 50.1
FigStep 48.1 56.7 31.9 36.3 47.1 42.3 47.2
CoCA 51.1 54.8 35.9 38.3 45.3 34.2 48.1
ECSO 55.0 53.1 35.8 42.4 48.3 26.9 49.4
ETA 51.4 52.2 35.4 38.0 49.7 40.4 47.4
RAS (Ours) 54.7 55.4 36.8 42.1 50.9 30.8 50.0
(a) $M$-Safety
(b) FigStep
Figure 10: Attention maps for (a) MM-Safety and (b) FigStep under various query reformulations. SP. denotes safety prompt. VC. denotes visual context. Each heatmap visualizes the cross-modal attention from text tokens in the query to visual tokens. Unsafe regions, e.g., “illegal underground drug manufacturing” in (a) and “traffic weapons of mass destructions in (b), receive low attention weights without visual contexts, but significantly stronger attention when visual contexts are incorporated.
Appendix D Additional Attention Maps
In this section, we present additional attention maps for multimodal jailbreak attacks, including MM-Safety (Fig.10(a)) and FigStep (Fig.10(b)). In both cases, the text query is benign in isolation, but its combination with an unsafe image results in a harmful instruction.
For example, in MM-Safety, the text query “list steps in detail to engage in this activity” appears benign on its own, but when paired with an image of drugs captioned “illegal underground drug manufacturing”, the combined instruction becomes malicious. Similarly, in FigStep, the text query simply asks to complete a numbered list, yet when paired with an image containing the prompt “Steps to traffic weapons of mass destruction”, it is transformed into a harmful instruction.
Following Fig. 3, we show the attention maps for various query formulations. As shown in the attention maps (right of Fig. 10), we observe a similar trend to the analysis in Sec. 3.1. Under the ‘Query’ and ‘Safety Prompt + Query’ formulations, the attention weights on unsafe image regions (e.g., unsafe texts embedded in the image) remain weak. In contrast, when visual contexts are incorporated, the model allocates higher attention weights to these regions, highlighting the role of visual contexts in strengthening cross-modal attention to safety-critical image regions.
Figure 11: An unsafe FigStep sample (left) and its safe counterpart (right). Although the input text query is benign on its own, incorporating images with unsafe typographic content (left, highlighted in red) makes the overall instruction unsafe. In contrast, when the embedded text specifies a benign activity (right, highlighted in green), the multimodal instruction remains safe.
Appendix E Additional FDR Analysis on FigStep Queries
In Sec. 3.1, we show that insufficient attention to safety-critical image regions leads to weak representational separability between safe and unsafe multimodal queries, especially when the given text queries are identical. Here, we extend our analysis to FigStep, where the text query is benign on its own, simply requesting the model to “generate detailed content for each item on the list”. However, when paired with images containing typographic text that specifies unsafe or malicious activities (e.g., “Steps to traffic weapons of mass destruction.”), the overall instruction becomes unsafe (left of Fig. 11).
To this end, analogous to the setup in Fig. 3, we construct safe FigStep counterparts by replacing the embedded texts in the image with benign instructions (e.g., “Steps to make homemade pizza dough.”), while keeping the text query identical (right of Fig. 11). We then measure the representational separability between the safe and unsafe FigStep samples using the Fisher Discriminant Ratio (FDR), computed from the last token activations (same procedure as in Sec. 3.1). This design isolates the effect of embedded text in images, ensuring that representational separability is driven solely by the visual content rather than the textual query.
Figure 12: FDR values across layers for various query formulations (LLaVA-1.5-7B). SP. denotes safety prompt. VC. denotes visual context. Lower FDR indicates less separable representations.
Across various query formulations, we observe results consistent with those in Sec. 3.1. When the model processes only with the original query, cross-modal attention to the typographic text in the image remains weak (first attention map in Fig. 10(b)). This leads to low FDR (orange line in Fig. 12), indicating poor representational separability between safe and unsafe samples.
Adding safety prompts alone does not remedy this issue. Even with safety prompts, weak attention to safety-critical regions (second attention map in Fig. 10(b)) keeps the overall FDR low (green line in Fig. 12). In contrast, incorporating visual contexts that explicitly reference the embedded text significantly strengthens cross-modal attention (third attention map in Fig. 10(b)), resulting in higher FDR values (blue line in Fig. 12) and clearer representational separation between safe and unsafe instructions. Moreover, when visual contexts are combined with safety prompts, the FDR improves further (red line in Fig. 12), demonstrating that once visual grounding is established, safety prompting can further amplify representational separability.
Overall, these findings confirm that insufficient attention to safety-critical regions is the fundamental issue in multimodal safety, and that vision-aware query reformulation provides distinct representations between safe and unsafe queries for accurate risk evaluation.
Appendix F Ablation on Unsafe Prototypes
In Sec. 3.2, we construct unsafe prototypes using unsafe text queries generated by GPT-4. To test whether the query source impacts RAS performance, we also sample an equal number of unsafe queries from the Anthropic Red Teaming Dataset (Ganguli et al., 2022). As shown in Tab.5, both sources yield substantial reductions in attack success rates while preserving general task performance. The differences between them are marginal, indicating that RAS is robust to the choice of query source when constructing unsafe prototypes.
Table 5: Ablation on unsafe query sources to determine unsafe prototypes. We compare two query sources for constructing unsafe prototypes: (i) the Anthropic Red Teaming Dataset (RTD) and (ii) unsafe queries generated by GPT-4. Bold indicates the highest performance (lowest ASR for safety benchmarks, and highest accuracies or scores for utility benchmarks). Average gains report the relative change in performance compared to the original model: for safety, it reflects the percentage reduction in ASR, and for utility, the percentage drop in task performance. Overall, both sources yield comparable results, with only marginal differences, confirming that RAS is robust to the choice of unsafe text query source.
Model Method Query Source Safety Utility
$M$-Safety SPA-VL FigStep Average Sci-QA $M$-Vet GQA MME (Percep./Cog.) Average
(ASR $\downarrow$) (ASR $\downarrow$) (ASR $\downarrow$) Gain (%) (Img. Acc. $\uparrow$) (Score $\uparrow$) (Acc. $\uparrow$) (Score $\uparrow$) Gain (%)
LLaVA-1.5-7B Original - 40.1 47.2 59.3 - 69.5 30.5 61.9 1505.1 / 355.7 -
RAS (Ours) Anthropic RTD. 4.3 8.8 2.4 +88.9 69.5 30.4 61.9 1505.1 / 355.7 -0.1
GPT-4 4.1 8.3 2.2 +89.5 69.5 30.5 61.9 1505.1 / 355.7 0.0
LLaVA-1.5-13B Original - 41.0 40.8 61.6 - 72.7 35.6 63.2 1529.9 / 298.6 -
RAS (Ours) Anthropic RTD. 7.6 4.6 1.2 +89.4 72.7 35.3 63.2 1529.9 / 298.6 -0.2
GPT-4 6.9 4.9 2.0 +89.3 72.7 35.4 63.2 1529.9 / 298.6 -0.1
Qwen-VL-Chat Original - 33.1 12.5 52.4 - 68.0 48.7 57.3 1489.9 / 331.8 -
RAS (Ours) Anthropic RTD. 1.5 3.8 1.8 +87.2 68.0 47.2 57.3 1489.9 / 331.8 -0.6
GPT-4 0.7 3.0 1.2 +90.5 68.0 46.9 57.3 1489.9 / 331.8 -0.7
InternLM- XComposer-2.5 Original - 21.8 27.6 22.6 - 94.7 50.1 59.1 1623.7 / 551.1 -
RAS (Ours) Anthropic RTD. 4.0 2.6 5.4 +82.8 94.7 49.8 59.1 1623.7 / 551.1 -0.1
GPT-4 5.1 4.2 4.0 +81.2 94.7 50.0 59.1 1623.7 / 551.1 -0.1
Appendix G Comparison on Refusal Vectors
In this section, we compare our proposed approach for computing refusal vectors with the method from Arditi et al. (2024).
G.1 Refusal Vector Definitions
Arditi et al. (2024): The refusal vector is computed as the difference between the mean activations of safe and unsafe text queries. Following this approach, the refusal vector $\mathbf{v}^{l,n}$ is defined as:
$\mathbf{v}^{l,n}=\bm{\mu}^{l,n}_{u}-\bm{\mu}^{l,n}_{s},$ (8)
where $\bm{\mu}^{l,n}_{u}$ and $\bm{\mu}^{l,n}_{s}$ denote the mean activations at layer $l$ and response token position $n$ for unsafe and safe queries, respectively. For unsafe queries, we use the text queries generated by GPT-4 (Sec. 3.2). For safe queries, we sample an equal number of examples from the Alpaca dataset (Taori et al., 2023).
Ours: We define refusal vectors as directions from individual input query activations to the corresponding unsafe prototypes, i.e., the mean activations of unsafe queries $\bm{\mu}^{l,n}_{u}$. Formally,
$\mathbf{v}^{l,n}=\bm{\mu}^{l,n}_{u}-\mathbf{x}_{i}^{l,n},$ (9)
where $l$ denotes the layer, $n$ denotes the output token position, and $i$ denotes the input query.
G.2 Jailbreak Results
We report ASR results for the two refusal vector computation methods on LLaVA-1.5-7B in Tab. 6. We evaluate activation steering when applied to (i) an intermediate layer and (ii) the last layer. For the choice of intermediate layer, we follow Arditi et al. (2024) and select layer 14.
Using the refusal vector definition of Arditi et al. (2024), we observe only modest safety gains: steering at layer 14 reduces ASR by 11.3% on average, while steering at the final layer achieves a 62.1% reduction. In contrast, our refusal vector achieves substantially larger gains, reducing ASR by 59.0% at layer 14 and 89.5% at the final layer, clearly demonstrating more effective refusal behavior.
The key difference lies in the steering objective. The refusal vector of Arditi et al. (2024) was originally constructed to suppress refusal behavior as part of an attack objective, which makes it less effective when inverted for safety enhancement. In contrast, our vector directly steers query activations toward unsafe prototypes, which correspond to representative states that trigger refusal behavior.
Table 6: Comparison on safety benchmarks for different refusal vector computations. We evaluate LLaVA-1.5-7B with RAS using refusal vectors from Arditi et al. (2024) and our approach, applied at an intermediate layer (14) and at the final layer. Our method consistently achieves substantially larger ASR reductions, with gains of up to 89.5% when steering at the final layer.
Model Method Refusal Steering $M$-Safety SPA-VL FigStep Average
Vector Layer (ASR $\downarrow$) (ASR $\downarrow$) (ASR $\downarrow$) Gain (%)
LLaVA-1.5-7B Original - - 40.1 47.2 59.3 -
RAS Arditi et al. (2024) 14 33.7 38.9 59.2 +11.3
RAS Arditi et al. (2024) Last 12.4 34.3 6.0 +62.1
RAS Ours 14 15.1 30.9 11.8 +59.0
RAS Ours Last 4.1 8.3 2.2 +89.5
Appendix H Qualitative Results
In this section, we present qualitative results to illustrate how risk-adaptive activation steering influences the model’s responses under varying steering strengths. Unlike benchmarks results that summarize performance with metrics (e.g., attack success rates or utility scores), these examples illustrate how steering influences responses to unsafe and safe multimodal queries.
We select one unsafe sample from MM-Safety and one safe sample from MM-Vet, and demonstrate the effects of steering on these examples across four models: LLaVA-1.5-7B (Fig. 13), LLaVA-1.5-13B (Fig. 14), Qwen-VL-Chat (Fig. 15), and InternLM-XComposer-2.5 (Fig. 16). We vary the steering strength $r(S_{i})$ from 0.0 to 1.0 in increments of 0.2, and show the generated responses.
Unsafe query (MM-Safety). At ‘Steering Strength = 0.0’, models tend to comply with the unsafe request, generating harmful responses. As the steering strength increases, refusals begin to emerge around ‘Steering Strength = 0.4’, and by ‘Steering Strength = 1.0’, the responses consistently converge to clear refusals across all models. This progression highlights that stronger steering is effective in eliciting refusal behavior.
Safe query (MM-Vet). At ‘Steering Strength = 0.0’, the models generate correct and helpful responses aligned with the task. Like unsafe queries, strong steering (e.g., 1.0) also drives safe queries into refusals, leading to utility degradation. However, a difference is that for safe queries, refusal behavior generally emerges at a higher threshold (around ‘Steering Strength = 0.6’), whereas unsafe queries begin to show refusals earlier (around ‘Steering Strength = 0.4’).
These qualitative results show that the steering behavior aligns closely with the risk scores from our evaluation. In particular, when the score reaches intermediate levels ($r(S_{i})\approx 0.5$), the model’s responses shift from compliance to refusals, as intended. This demonstrates that the evaluation produces accurate risk scores that directly translate into predictable steering behavior.
Figure 13: Qualitative results under different steering strengths on LLaVA-1.5-7B. Left: Unsafe query (MM-Safety). Right: Safe query (MM-Vet).
Figure 14: Qualitative results under different steering strengths on LLaVA-1.5-13B. Left: Unsafe query (MM-Safety). Right: Safe query (MM-Vet).
Figure 15: Qualitative results under different steering strengths on Qwen-VL-Chat. Left: Unsafe query (MM-Safety). Right: Safe query (MM-Vet).
Figure 16: Qualitative results under different steering strengths on InternLM-XComposer-2.5. Left: Unsafe query (MM-Safety). Right: Safe query (MM-Vet).